# Establishing a Foundation for Tetun Text Ad-Hoc Retrieval: Stemming, Indexing, Retrieval, and Ranking

### v2

@@ 9.5. Short-Text Retrieval Results
For short-text retrieval, we indexed document titles and then used them for retrieval. The impact of each preprocessing technique on retrieval effectiveness for short-text retrieval is presented Table 23. The scores highlighted in red indicate values lower than the baseline. As observed, removing accents and applying all stemmer variants did not improve retrieval effectiveness compared to the baseline. Stopword removal shows inconsistent results across models and metrics; it generally performs better only at top-20 cutoffs and in overall MAP and nDCG, while demonstrating poor performance at top-5 and 10 cutoffs. When focusing solely on the impact of stemming on retrieval effectiveness, the light stemmer variant slightly outperforms the moderate and heavy variants across all cutoffs, with moderate and heavy variants performing similarly.
@@
Based on the preliminary results in Table 23, we tested all possible combinations of techniques that outperformed the baseline, including stopword removal and the light stemmer variant, for further comparison. The best results from these combined preprocessing techniques for indexing document titles are presented in Table 24. The findings reveal that removing apostrophes and hyphens significantly improves the retrieval effectiveness compared to the baseline.
Among the retrieval models, DFR BM25 consistently achieved the highest performance across all metrics up to top-10 cutoffs, with notable scores in Precision@5 (0.8881), Precision@10 (0.8390), MAP@5 (0.1589, with stopword removal), MAP@10 (0.2804), nDCG@5 (0.7512), and nDCG@10 (0.7356). Specifically, DFR BM25 demonstrated improvements of up to 30.35% over the baseline and up to 8.19% over individual text preprocessing techniques on multiple metrics, and modest gains of up to 5.54% in MAP@5 over other retrieval models within the same setting. DFR BM25 achieved the same Precision@5 as BM25 and MAP@5 as TF-IDF while showing slightly higher scores in the other cutoffs.
Table 24. Effectiveness of Combined Text Preprocessing Techniques in Short-Text Retrieval.
@@
The removal of hyphens and stopwords also proved beneficial, particularly with Hiemstra LM, which recorded the highest scores for Precision@20 (0.7305), MAP@20 (0.4372), nDCG@20 (0.7152), MAP (0.7040), and nDCG (0.8289, with apostrophe removal). This combination improves retrieval effectiveness overall, with Hiemstra LM consistently showing gains at top-20 cutoffs as well as global MAP and nDCG, demonstrating improvements of up to 15.60% over the baseline and up to 8.33% over individual text preprocessing techniques in multiple metrics, with marginal gains of up to 3.94% over other retrieval models within the same setting.
Stemming, on the other hand, continued to show minimal impact on retrieval effectiveness, even when combined with other text preprocessing techniques that outperformed the baseline. The best result among all combinations with the light stemmer variant involved the removal of hyphens and stopwords. Although this combination produced competitive results, it did not surpass the strategy that excluded stemming.
@@ 9.6. Long-Text Retrieval Results
For long-text retrieval, we indexed the document content and performed retrieval using this index. The impact of each preprocessing technique on retrieval effectiveness is presented in Table 25. The results reveal different impacts compared to short-text retrieval, with apostrophe and accent removal showing no specific benefit for BM25-based models in Precision and MAP compared to the baseline. Accent removal shows no improvement in retrieval effectiveness at P@10, P@20, and MAP metrics with the Dirichlet LM model. However, there is an indication of improvement with Hiemstra LM at the MAP and nDCG cutoffs, as well as overall nDCG. In contrast, hyphen removal, stopword removal, and stemming demonstrate clear benefits for long-text retrieval. Among the stemming variants, the light stemming continues to outperform the moderate and heavy variants, though the margin over the moderate variant is minimal.
@@
The results in Table 25 were further analyzed by combining text preprocessing techniques to better understand their impact on long-text retrieval effectiveness. Combined these preprocessing techniques produced varied impacts across evaluation metrics and models. Hiemstra LM consistently demonstrated the best performance in most combinations and cutoffs, with the exception at P@5 and nDCG@5, where TF-IDF showed a slight performance advantage. Notable results with Hiemstra LM include the combination of apostrophe and hyphen removal, which achieved the highest scores at P@10 (0.5576), P@20 (0.4907), MAP@20 (0.2523), and nDCG@20 (0.4790).
Additional configurations, such as the inclusion of moderate stemming yielded the best results at MAP@5 (0.1645), while including stopword removal improved nDCG@10 (0.4760) and MAP (0.4358). Furthermore, Hiemstra LM exhibited effectiveness gains by combining hyphen and accent removal, particularly at MAP@10 (0.1645) and nDCG (0.6855). Hyphen removal alone continued to demonstrate its effectiveness in long-text retrieval, achieving the same highest score as the combination with apostrophe removal at P@20 (0.4907).
Table 26. Effectiveness of Combined Text Preprocessing Techniques in Long-Text Retrieval.
@@
In contrast, individual removal techniques, such as accent and apostrophe removal, generally underperformed when isolated, contributing minimally across different metrics (see Table 25). However, combining these techniques with hyphen and stopword removal along with moderate stemming resulted in strong performance, particularly with Hiemstra LM across most metrics and cutoffs. Although the heavy stemming variant slightly outperformed the moderate variant in individual preprocessing, the moderate variant gained a slight advantage when combined with other text preprocessing techniques.
@@
Hiemstra LM consistently achieved the highest scores with the combined preprocessing approach across all metrics, except at P@5, where TF-IDF slightly outperformed it. Hiemstra LM demonstrated average performance gains over the baseline, reaching up to 13.75% for Precision cutoffs, 18.01% for MAP cutoffs, 14.48% for nDCG cutoffs, 19.23% for MAP, and 14.44% for nDCG. Additionally, Hiemstra LM improved an average of up to 16.40% over the other models.
@@
In contrast, BM25 and DFR BM25 did not yield the best scores with any preprocessing strategies for long-text retrieval, suggesting that BM25-based models may be less suitable for long-document contexts. These findings highlight that text preprocessing techniques and retrieval models affect short- and long-text retrieval differently in Tetun.
9.7. Discussions
@@
In Tetun text ad-hoc retrieval, text preprocessing techniques that involve the removal of apostrophes and hyphens generally lead to performance gains of up to 8.19% with DFR BM25 in short-text retrieval up to the top-10 cutoffs. When stopword removal is combined with apostrophe and hyphen removal, additional gains are observed, particularly at MAP@5 with both DFR BM25 and TF-IDF models, and at nDCG with Hiemstra LM. When apostrophes are retained and only hyphens and stopwords are removed, Hiemstra LM shows improvements of up to 8.33% compared to other models at P@20, MAP@20, nDCG@20, and overall MAP. These results suggest that in short-text retrieval, targeting hyphen removal can enhance retrieval effectiveness.
@@
Furthermore, different strategies for handling stopwords and apostrophes yield varied benefits across models: retaining stopwords and removing apostrophes works better with DFR BM25 while retaining apostrophes and removing stopwords proves more effective with Hiemstra LM at higher cutoffs and overall. Accent removal, however, does not improve retrieval effectiveness in short-text retrieval, regardless of whether it is applied individually or in combination with other text preprocessing techniques.
Additionally, removing stopwords does not lead to noticeable improvements in retrieval effectiveness, for example, at P@20, the gain is only 0.58% improvement compared to Hiemstra LM with the removal of apostrophes and hyphens. Similarly, MAP@20 shows a marginal improvement of 0.76%, nDCG@20 a gain of 0.07%, and overall MAP an improvement of 1.22% compared to Hiemstra LM with the removal of apostrophes and hyphens.
@@
Our findings on the impact of stopword removal align with those reported by Savoy (Savoy, 1999) in experiments with French, where retaining stopwords yielded the best results with the BM25 model. This outcome is also consistent with Ghosh and Bhattacharya (Ghosh and Bhattacharya, 2017), who observed variability in retrieval effectiveness within the same language and found no significant improvement with stopword removal. Moreover, our results are similar to those of Dolamic and Savoy (Dolamic and Savoy, 2010), who reported that retaining stopwords improved retrieval effectiveness in Hindi, with no significant difference in effectiveness for Marathi and Bengali.
@@
With respect to stemming, it does not improve retrieval effectiveness in Tetun, whether applied independently or in combination with other preprocessing techniques. This finding is consistent with our preliminary experiments on Tetun text retrieval reported earlier (de Jesus, 2022) and with the results of Flores and Moreira (Flores and Moreira, 2016), who observed similar results in their experiments with Dutch, Italian, Spanish, and English.
@@
In contrast, several studies on LRLs, including those by Sahu and Paul (Sahu and Pal, 2023), Sahu et al. (Sahu et al., 2023), and Adriani et al. (Adriani et al., 2007), reported positive impacts of stemming on retrieval effectiveness. This difference is likely due to the language-specific characteristics of Tetun, which may respond differently to stemming compared to other languages. Moreover, although intrinsic evaluations suggested that moderate and heavy stemmers were slightly more accurate than light stemmer, this did not consistently translate into improved retrieval performance. This finding aligns with Flores and Moreira’s (Flores and Moreira, 2016) observation that higher stemmer accuracy does not always lead to improvement of retrieval effectiveness.
@@
For long-text retrieval, Hiemstra LM consistently demonstrated superior performance across most metrics and cutoffs. Hyphen removal yielded performance gains across different models and metrics, particularly when combined with apostrophe removal, achieving high scores at P@10, P@20, MAP@20, and nDCG@20. Adding moderate stemming to this combination resulted in the highest MAP@5 score, although the difference was minimal compared to strategies without stemming.
@@
The removal of apostrophes showed inconsistent results as an individual preprocessing technique but was more effective when combined with the removal of hyphens. Similarly, accent removal had inconsistent impacts across models and metrics when applied alone; however, in combination with other techniques, it produced a 16.40% improvement in MAP@10 over other models. This outcome aligns with the findings of Savoy (Savoy, 1999) in his study on French, where accent removal slightly improved retrieval effectiveness in long-text retrieval but decreased effectiveness in short-text retrieval for several retrieval strategies.
9.8. Conclusions
@@
Short-text retrieval is more effective than long-text retrieval for Tetun text ad-hoc retrieval. Splitting compound words by removing hyphens in both queries and documents, combined with apostrophe removal, improves retrieval efficiency and effectiveness. The removal of stopwords shows no significant impact while retaining accents is crucial to effective text retrieval in Tetun.
@@
Stemming does not improve retrieval effectiveness in short-text retrieval, while a moderate variant combined with other preprocessing techniques shows a slight improvement in long-text retrieval at MAP@5, the score remains lower than 0.1. This limited impact suggests that morphological reduction through stemming may not be essential for effective text retrieval in Tetun, likely due to the relatively simple morphology of the language with few inflectional affixes.
@@
Overall, the experimental results show that text preprocessing, including the removal of hyphens and apostrophes, significantly enhances the effectiveness of Tetun text ad-hoc retrieval. DFR-BM25 performs best up to top-10 cutoffs but appears to be more sensitive to stopword removal, while Hiemstra LM delivers the most significant benefits beyond top-20 cutoffs and overall MAP and nDCG.
@@ 10. Conclusions and Future Work
This study details the development of Tetun text retrieval and establishes the first baselines for the ad-hoc retrieval task. To support Tetun text retrieval, three essential resources were developed: Labadain-Stopwords, Labadain-Stemmer, and Labadain-Avaliadór. These resources will be made publicly available to future IR and NLP researchers.
@@
The Labadain-Stopwords includes 160 Tetun stopwords, a total comparable to those reported for other low-resource languages (LRLs) in the literature. The Labadain-Stemmer is available in three variants—lightmoderate, and heavy—with the light variant specifically targeting Portuguese loanwords commonly used in Tetun. Due to the prominence of these loanwords, the Portuguese stemmer from Snowball was adapted for Tetun by aligning it with the linguistic characteristics of Portuguese-derived words in Tetun. This adaptation method can be extended to other LRLs with a similar linguistic nature.
@@
The Labadain-Avaliadór consists of 59 topics, 33,550 documents, and 5,900 qrels, with an average of 36.76 relevant documents per query. This average reflects a balanced representation of the relevant documents in the collection. The balance arises from the fact that the queries were derived from real-world search logs sourced from two platforms: Google Search Console logs for Timor News and logs from searches performed using the search functionality of Timor News. This ensures that queries and documents reflect real-world scenarios where the documents actually exist in the Labadain-30k+ dataset (de Jesus and Nunes, 2024b) used as the document collection.
@@
Our investigation involved experimenting with different retrieval strategies, with a focus on the impact of various text preprocessing techniques tailored based on the linguistic characteristics of Tetun. We initially hypothesized that text preprocessing techniques would improve the effectiveness of Tetun text retrieval, and our findings confirmed this hypothesis. In response to our research questions, we conclude that removing hyphens to split compound words into individual words, combined with apostrophe removal, enhances retrieval effectiveness overall in both short- and long-text retrieval, with short-text retrieval being the most effective approach for Tetun.
@@
For the retrieval and ranking models, DFR BM25 performs effectively with short-text retrieval up to the top-10 cutoff but shows slightly lower effectiveness when stopwords are removed. Meanwhile, Hiemstra LM consistently demonstrates effective performance across various metrics, particularly beyond the top-10 cutoffs and in overall MAP and nDCG. These findings suggest that Hiemstra LM is more effective for Tetun text retrieval when retrieving more than ten documents is prioritized.
The effectiveness of Hiemstra LM for Tetun text retrieval is consistent with the findings of Sahu and Pal (Sahu and Pal, 2023) in their study on Sanskrit, a script-based language spoken in India. Although Tetun uses a Latin script, both languages demonstrate comparable retrieval effectiveness with the Hiemstra LM model in short-text retrieval measured by MAP. This suggests that Hiemstra LM may adapt effectively to LRLs, regardless of their linguistic or script characteristics.
@@
Future work will investigate semantic search, which captures the contextual meaning of queries and documents rather than relying solely on exact term matches. Integrating large language models (LLMs) into retrieval tasks may open new avenues for enhancing retrieval effectiveness and better align retrieval system with user information needs. Additionally, investigating user search behavior influenced by LLM advancements may reveal evolving patterns in user search intent and information needs, particularly in LRLs like Tetun. Insights from these trends could guide the design of retrieval systems that effectively adapt to changing search behaviors.

### v6

@@ 9.5. Short-Text Retrieval Results
For short-text retrieval, document titles were indexed and used for retrieval. The impact of each preprocessing technique on retrieval effectiveness is presented Table 24. Scores highlighted in red indicate values lower than the baseline. As observed, removing accents and applying all stemmer variants did not improve retrieval effectiveness compared to the baseline. Stopword removal yielded inconsistent results across models and metrics. It generally demonstrated improved performance only at the top-20 cutoffs (P@20, MAP@20, and NDCG@20) and overall MAP and NDCG, while exhibiting lower performance at top-5 and top-10 cutoffs across all metrics. Regarding the impact of stemming on retrieval effectiveness, the light stemmer variant slightly outperformed the moderate and heavy variants across all cutoffs, with the moderate and heavy variants performing similarly.
@@
Based on the preliminary results in Table 24, all combinations of preprocessing techniques that outperformed the baseline, including stopword removal and the light stemmer variant, were selected for further comparison. The best results from these combined preprocessing techniques are presented in Table 25. The findings indicate that removing apostrophes and hyphens significantly enhances retrieval performance compared to the baseline.
Among the retrieval models, DFR BM25 consistently delivered the highest performance across all metrics up to top-10 cutoffs. Notable scores include P@5 (0.8881), P@10 (0.8390), MAP@5 (0.1589, removing stopwords), MAP@10 (0.2804), NDCG@5 (0.7512), and NDCG@10 (0.7356). DFR BM25 demonstrated relative improvements of up to 30.35% over the baseline and up to 8.19% over individual preprocessing techniques on multiple metrics, along with modest relative gains of up to 5.54% in MAP@5 over other retrieval models within the same settings. While DFR BM25 achieved the same P@5 as BM25 and MAP@5 as TF-IDF, it demonstrated slightly higher scores in other cutoffs.
Table 25. Effectiveness of Combined Text Preprocessing Techniques in Short-Text Retrieval. Values highlighted with a green background indicate the best score at the respective metric and cutoff.
@@
The removal of hyphens and stopwords also proved beneficial, particularly with Hiemstra LM, which demonstrated the highest scores for P@20 (0.7305), MAP@20 (0.4372), NDCG@20 (0.7152), and overall MAP (0.7040) and NDCG (0.8289, with apostrophe removal). This combination enhanced retrieval effectiveness overall, with Hiemstra LM consistently showing relative improvements at top-20 cutoffs and in global MAP and NDCG. It demonstrated relative improvements of up to 15.60% over the baseline and up to 8.33% over individual text preprocessing techniques in multiple metrics, with marginal relative gains of up to 3.94% over other retrieval models within the same setting.
Stemming, however, had minimal impact on retrieval effectiveness, even when combined with other text preprocessing techniques that outperformed the baseline. The best result involving the light stemmer variant was achieved when paired with hyphen and stopword removal. While this combination produced competitive results, it did not surpass the performance of strategies that excluded stemming.
@@ 9.6. Long-Text Retrieval Results
Document content was indexed and used for long-text retrieval. The impact of each preprocessing technique on retrieval effectiveness is summarized in Table 26. The results differ from those of short-text retrieval: apostrophe and accent removal did not provide any measurable benefit for BM25-based models in terms of Precision or MAP relative to the baseline. Likewise, accent removal offered no improvement in P@10, P@20, or MAP with the Dirichlet LM model.
However, there was evidence of improvement with Hiemstra LM in MAP and NDCG at various cutoffs and in overall NDCG. In contrast, hyphen removal, stopword removal, and stemming demonstrated clear advantages for long-text retrieval. Among the stemming variants, the light stemmer consistently outperformed the moderate and heavy variants, though the margin of improvement over the moderate variant was minimal.
@@
The results presented in Table 26 were further analyzed by combining text preprocessing techniques to better understand their collective impact on long-text retrieval effectiveness. These combined preprocessing techniques yielded varied impacts across evaluation metrics and models. Hiemstra LM consistently outperformed other models in most combinations and cutoffs, except in P@5 and NDCG@5, where TF-IDF exhibited a slight performance advantage. Notable results for Hiemstra LM include the combination of apostrophe and hyphen removal, which achieved the highest scores at P@10 (0.5576), P@20 (0.4907), MAP@20 (0.2523), and NDCG@20 (0.4790).
Additional configurations, such as incorporating moderate stemming, yielded the best results at MAP@5 (0.1645), while including stopword removal improved NDCG@10 (0.4760) and overall MAP (0.4358). Furthermore, Hiemstra LM showed notable gains in effectiveness when combining hyphen and accent removal, particularly at MAP@10 (0.1645) and NDCG (0.6855). Hyphen removal alone continued to demonstrate its effectiveness in long-text retrieval, achieving the same highest score as its combination with apostrophe removal at P@20 (0.4907).
Table 27. Effectiveness of Combined Text Preprocessing Techniques in Long-Text Retrieval. Values highlighted with a green background indicate the best score at the respective metric and cutoff.
@@
In contrast, individual preprocessing techniques, such as accent and apostrophe removal, generally underperformed when applied in isolation, contributing minimally to retrieval effectiveness across different metrics (see Table 26). However, combining these techniques with hyphen and stopword removal, along with moderate stemming, resulted in a strong performance, particularly with Hiemstra LM across most metrics and cutoffs. While the heavy stemming variant slightly outperformed the moderate variant when applied individually, the moderate variant showed a slight advantage when combined with other preprocessing techniques.
@@
Hiemstra LM consistently delivered the highest scores with the combined preprocessing approach across all metrics, except at P@5, where TF-IDF marginally outperformed it. Hiemstra LM demonstrated notable performance improvements over the baseline, achieving gains of up to 13.75% for Precision cutoffs, 18.01% for MAP cutoffs, 14.48% for NDCG cutoffs, 19.23% for overall MAP, and 14.44% for overall NDCG. Additionally, Hiemstra LM achieved an average relative improvement of up to 16.40% over the other retrieval models.
@@
Conversely, BM25 and DFR BM25 did not produce the best results with any preprocessing strategies for long-text retrieval, suggesting that BM25-based models may be less effective in handling long-document contexts. These findings underscore the different impacts of text preprocessing techniques and retrieval models on short- and long-text retrieval tasks in Tetun.
9.7. Discussion
@@
In Tetun ad-hoc text retrieval, text preprocessing techniques that involve the removal of apostrophes and hyphens generally lead to performance gains of up to 8.19% with DFR BM25 in short-text retrieval up to the top-10 cutoffs. When stopword removal is combined with apostrophe and hyphen removal, further gains are observed, particularly at MAP@5 with both DFR BM25 and TF-IDF models and at NDCG with Hiemstra LM. When apostrophes are retained and only hyphens and stopwords are removed, Hiemstra LM shows improvements of up to 8.33% compared to other models at P@20, MAP@20, NDCG@20, and overall MAP. These results suggest that in short-text retrieval, targeting hyphen removal can enhance retrieval effectiveness.
@@
Different strategies for handling stopwords and apostrophes yield varied benefits across retrieval models. Retaining stopwords while removing apostrophes proves more effective with DFR BM25, whereas retaining apostrophes while removing stopwords shows better results with Hiemstra LM at higher cutoffs and overall. Accent removal, however, does not enhance retrieval effectiveness in short-text retrieval, whether applied individually or in combination with other preprocessing techniques.
@@
Removing stopwords does not lead to noticeable improvements in retrieval effectiveness. For instance, at P@20, the relative gain is limited to 0.58% compared to Hiemstra LM with apostrophe and hyphen removal. Similarly, MAP@20 improves by 0.76%, NDCG@20 by 0.07%, and overall MAP by 1.22% relative to the baselines. These findings align with Savoy’s (Savoy, 1999) experiments with French, which showed that retaining stopwords yielded better results with the BM25 model. Similar conclusions were reported by Ghosh and Bhattacharya (2017), who found variability in retrieval effectiveness within the same language and no significant improvement from stopword removal. Additionally, Dolamic and Savoy (2010) observed that retaining stopwords improved retrieval effectiveness in Hindi, with minimal differences for Marathi and Bengali.
@@
Regarding stemming, it does not improve retrieval effectiveness in Tetun, whether applied independently or in combination with other preprocessing techniques. This result is consistent with findings from preliminary experiments on Tetun text retrieval (de Jesus, 2022) and Flores and Moreira’s (Flores and Moreira, 2016) experiments with Dutch, Italian, Spanish, and English. However, studies on other LRLs, including those by Sahu and Pal (2023) (Sanskrit), Sahu et al. (2023) (Urdu), and Adriani et al. (2007) (Indonesian), reported positive impacts of stemming on retrieval effectiveness. This discrepancy likely arises from Tetun’s language-specific characteristics, which may respond differently to stemming than other languages.
@@
While intrinsic evaluations suggested that the moderate and heavy stemmers were slightly more accurate than the light stemmer, these differences did not consistently translate into improved retrieval performance. These findings align with Flores and Moreira’s (Flores and Moreira, 2016) results that higher stemmer accuracy does not always improve retrieval effectiveness.
@@
For long-text retrieval, Hiemstra LM consistently outperformed other models across most metrics and cutoffs. Hyphen removal delivered performance gains across various metrics and models, particularly when combined with apostrophe removal, achieving high scores at P@10, P@20, MAP@20, and NDCG@20. Adding moderate stemming to this combination yielded the highest MAP@5 score, though the improvement was minimal compared to strategies without stemming.
@@
Apostrophe removal showed inconsistent results as an individual preprocessing technique but became more effective when combined with hyphen removal. Similarly, accent removal had varying impacts across models and metrics when applied alone, but combined with other techniques, it achieved a relative improvement of 16.40% in MAP@10 over other models. This outcome aligns with Savoy’s (Savoy, 1999) findings on French, where accent removal slightly improved retrieval effectiveness in long-text retrieval but negatively affected short-text retrieval for several strategies.
9.8. Conclusion
@@
The most effective retrieval strategy for Tetun ad-hoc text retrieval involves short-text retrieval (i.e., using document titles) combined with targeted preprocessing techniques, specifically splitting compound words by removing hyphens and eliminating apostrophes from queries and documents. Together, these techniques enhance both retrieval efficiency and effectiveness. The removal of stopwords shows minimal impact, while preserving accents proves crucial for effective text retrieval in Tetun, likely due to the language’s reliance on diacritics to distinguish word meanings.
@@
Stemming does not improve retrieval effectiveness in short-text retrieval. In long-text retrieval, the moderate stemming variant, when combined with other preprocessing techniques, shows slight improvements at MAP@5; however, the gains remain marginal and are still lower than the best-performing score achieved in short-text retrieval, representing -39.25% relative performance drop. This minimal impact suggests that morphological normalization through stemming is not essential for effective text retrieval in Tetun, likely due to the language’s relatively simple morphology and minimal use of inflectional affixes.
@@
The experimental results show that text preprocessing techniques, particularly the removal of hyphens and apostrophes, are crucial in enhancing the effectiveness of Tetun ad-hoc text retrieval. Among the retrieval models, DFR BM25 performs best for cutoffs up to the top-10 but exhibits greater sensitivity to stopword removal. In contrast, Hiemstra LM provides the most substantial improvements at top-20 cutoffs and beyond, particularly in overall MAP and NDCG.
Overall, this study demonstrates the importance of thoroughly investigating language-specific preprocessing strategies by segmenting them into distinct stages and systematically integrating them to uncover the linguistic components that impact retrieval effectiveness (see Figure 13). It emphasizes the benefit of developing approaches tailored to the unique characteristics of each language, rather than relying solely on established techniques in the literature. The study introduces a detailed and adaptable methodology for establishing ad-hoc text retrieval baselines, particularly for LRLs where such benchmarks have not yet been established. By accounting for the unique morphological and syntactic features of each language, researchers can design more effective and linguistically appropriate retrieval strategies.
@@ 10. Conclusions and Future Work
This study presents the development of Tetun text retrieval and establishes the first baselines for the ad-hoc retrieval task. As part of this effort, we created three essential resources: Labadain-Stopwords (de Jesus and Nunes, 2025c), Labadain-Stemmer (de Jesus and Nunes, 2025b), and Labadain-Avaliadór (de Jesus and Nunes, 2025a). These resources are publicly available to the IR and NLP research community under the Creative Commons Attribution-ShareAlike license.
@@
Labadain-Stopwords contains 160 Tetun stopwords, a total comparable to those reported for other low-resource languages (LRLs) in the literature. Labadain-Stemmer is available in three variants—lightmoderate, and heavy—with the light variant specifically targeting Portuguese loanwords commonly used in Tetun. Due to the prominence of these loanwords, the Portuguese stemmer from Snowball was adapted for Tetun by adapting it to the linguistic characteristics of Portuguese-derived words in Tetun. This method can be extended to other LRLs with a similar linguistic nature.
@@
Labadain-Avaliadór consists of 59 topics, 33,550 documents, and 5,900 relevance judgments (qrels), with an average of 36.76 relevant documents per query. This average reflects a balanced representation of the relevant documents in the collection. The balance arises from the fact that the queries were derived from real-world search logs sourced from two platforms: Google Search Console logs for Timor News and logs from searches performed using the Timor News search functionality. This ensures that queries and documents reflect real-world scenarios where the documents exist in the Labadain-30k+ dataset (de Jesus and Nunes, 2024b), which is used as the document collection.
@@
Our investigation involved experimenting with different retrieval strategies, with a focus on the impact of various text preprocessing techniques tailored based on the linguistic characteristics of Tetun. We initially hypothesized that text preprocessing techniques would improve the effectiveness of Tetun text retrieval, and our findings confirmed this hypothesis. In response to our research questions, we conclude that removing hyphens to split compound words into individual words, combined with the removal of apostrophes, enhances retrieval effectiveness overall in both short- and long-text retrieval, with short-text retrieval being the most effective approach for Tetun.
@@
For the retrieval and ranking models, DFR BM25 performs effectively with short-text retrieval up to the top-10 cutoff but shows slightly lower effectiveness when stopwords are removed. Meanwhile, Hiemstra LM consistently demonstrates effective performance across various metrics, particularly beyond the top-10 cutoffs and in overall MAP and NDCG. These findings suggest that Hiemstra LM is more effective for Tetun text retrieval when more than ten documents are prioritized for retrieval. The effectiveness of Hiemstra LM for Tetun text retrieval is consistent with the findings of Sahu and Pal (2023) in their study on Sanskrit, a script-based language spoken in India. Although Tetun uses a Latin script, both languages demonstrate comparable retrieval effectiveness with the Hiemstra LM model in short-text retrieval, measured by MAP. This suggests that Hiemstra LM may adapt effectively to LRLs, regardless of their linguistic or script characteristics.
@@
Future work will investigate semantic search, which captures the contextual meaning of queries and documents rather than relying solely on exact term matches. Integrating large language models (LLMs) into retrieval tasks may open new avenues for enhancing retrieval effectiveness and better aligning retrieval systems with user information needs. Additionally, investigating user search behavior influenced by LLM advancements may reveal evolving patterns in user search intent and information needs, particularly in LRLs like Tetun. Insights from these trends could guide the design of retrieval systems that effectively adapt to changing search behaviors.